Can you help me solve this difficult complicated problem?
Suppose you deposited $625 in an account with an annual interest rate of 4% compounded semi-annually. Use this information to answer parts a,b and c
A) Find the Equation that gives the amount of money in the account after t years.
B) Find the amount of money in the account after 12 years (round to the nearest penny).
C) If the Interest were Compounded Continuously, how much money would the account be worth after 5 years (round to the nearest penny).
Thank you so much, I greatly appreciate it !!

Answer by stanbon(75887)

Can you help me solve this difficult complicated problem?
Suppose you deposited $625 in an account with an annual interest rate of 4% compounded semi-annually.
Use this information to answer parts a,b and c
A) Find the Equation that gives the amount of money in the account after t years.
A(t) = P(1+(r/n))^(n*t)
r is the annual rate
n is the # of times you compound each year
t is the # of years you compound
-------------------------------------
B) Find the amount of money in the account after 12 years (round to the nearest penny).
A(12) = 625(1+(0.04/2))^(2*12) = $1005.27
--------------------------------
C) If the Interest were Compounded Continuously, how much money would the account be worth after 5 years (round to the nearest penny).
Formula: A(t) = P*e^(r*t)
A(12) = 625*e^(0.04*15) = $1138.82
-----
Cheers,
Stan H.
--------